OpenStudy (anonymous):

Help please? Find the 6th term of the sequence

OpenStudy (anonymous):

\[ t_{1} = -4 ,t _{n} = 5t _{n-1} .\]

OpenStudy (triciaal):

check your formula

OpenStudy (anonymous):

Ok so t1 = -4 and every other term is 5 times previous So t2 = -4 * 5 and t3 = -4 * 5^2 So t6 = -4 * 5^5 which is -12500

OpenStudy (triciaal):